The addition reaction between 1-propene and water.
1
Expert's answer
2018-01-31T05:07:07-0500
The addition reaction between 1-propen and water
According to the Markovnicoff’s rule
CH2 =CH-CH3 + H2O = CH3-CH(OH)-CH3 propanol-2